George can rake a lawn in 5 hours. George and Juan together can rake the lawn in 2 hours. How long would it take for Juan to rake the lawn by himself? Use the following equation to solve this problem where J represents the time it would take Juan to rake the lawn:
2/5+2/J=1

Answer by richwmiller(17219)

2/5+2/J=1
2/j=3/5
10=3j
10/3=j
3 1/3=j